# Stringlet Excitation Model of the Boson Peak

### v2

@@ Abstract
The boson peak (BP), a low-energy excess in the vibrational density of states over the Debye contribution, is often identified as a characteristic of amorphous solid materials. Despite decades of efforts, its microscopic origin still remains a mystery. Recently, it has been proposed, and corroborated with simulations, that the BP might stem from intrinsic localized modes involving one-dimensional (1D) string-like excitations (“stringlets”). We build on a theory originally proposed by Lund that describes the localized modes as 1D vibrating strings, but we specify the stringlet size distribution to be exponential, as observed in simulations. We provide an analytical prediction for the BP frequency $\omega_{BP}$ in the temperature regime well below the observed glass transition temperature $T_{g}$. The prediction involves no free parameters and accords quantitatively with prior simulation observations in 2D and 3D model glasses based on inverse power law potentials. The comparison of the string model to observations is more uncertain when compared to simulations of an Al-Sm metallic glass material at temperatures well above $T_{g}$. Nonetheless, our stringlet theory of the BP naturally reproduces the softening of the BP frequency upon heating and offers an analytical explanation for the experimentally observed scaling with the shear modulus in the glass state and changes in this scaling in simulations of glass-forming liquids. Finally, the theoretical analysis highlights the existence of a strong damping for the stringlet modes above $T_{g}$, which leads to a large low-frequency contribution to the 3D vibrational density of states, observed in both experiments and simulations.
@@ I Introduction
Amorphous solids exhibit a variety of characteristic anomalies in their vibrational, thermodynamic and transport properties, when compared to the more familiar case of ordered crystalline matter Zeller and Pohl (1971); Ramos (2022). Among the different, and apparently universal Ramos (2020), anomalous properties of glassy materials, the boson peak (BP) is probably the most commonly discussed and controversial. This peak refers to an excess in the vibrational density of states (VDOS) normalized by the phonon density of states of Debye theory, $g(\omega)/\omega^{d-1}$, where $d$ is the number of spatial dimensions and $\omega$ the frequency. The BP is ubiquitously observed in amorphous systems Alexander (1998) and recently this feature has also been observed in crystalline systems at finite temperature (e.g., Takabatake et al. (2014); Moratalla et al. (2019)), so it is unclear whether it arises from structural disorder Baggioli and Zaccone (2019a).
@@
Several theoretical models have been proposed to rationalize the BP anomaly in glasses Elliott (1992); Schirmacher (2006); Schirmacher, Ruocco, and Scopigno (2007); Schirmacher, Ruocco, and Mazzone (2015); Léonforte et al. (2006); Buchenau et al. (1991); Gurevich, Parshin, and Schober (2003); Parshin, Schober, and Gurevich (2007); Götze and Mayr (2000); Baggioli and Zaccone (2019a); Taraskin et al. (2001); Chumakov et al. (2011); Galperin, Karpov, and Kozub (1989); Klinger and Kosevich (2002); Grigera et al. (2003), but a final verdict has not yet been reached. A common idea is that the BP represents a signature for the emergence of additional vibrational modes coexisting, but distinct from phonons. These “excess modes” have been suggested to arise from both structural disorder or from the anharmonicity in intermolecular interactions inherent to the liquid state and heated crystals. These commonly observed vibrational modes are expected to be effectively localized as they involve a rather limited number of atoms or molecular segments in complex molecules Gurevich, Parshin, and Schober (2003); Parshin, Schober, and Gurevich (2007); Schober and Oligschleger (1996); Lerner and Bouchbinder (2022, 2023); Moriel, Lerner, and Bouchbinder (2023); Rainone et al. (2021); Vural and Leggett (2011); Schober, Oligschleger, and Laird (1993). At large enough wave-vector, they also interact with phonons, making them overdamped, i.e., reaching the Ioffe-Regel (IR) limit Shintani and Tanaka (2008); Beltukov, Kozub, and Parshin (2013); Rufflé et al. (2006), leading to hindered thermal transport because of the consequent strong phonon scattering.
@@
Despite numerous experimental and simulation studies, a fully predictive theoretical model explaining the existence of these modes and how they give rise to the observed boson peak remains elusive. Recently, it has been recognized that, on atomic scales, localized string-like excitations that involve reversible particle exchange motion in the form of linear polymeric structures exist in glass-forming liquids. These dynamic structures have been termed “stringletsZhang et al. (2021a). See Fig.1 for a visual representation from simulations. Moreover, it has been shown Zhang and Douglas (2013a); Zhang et al. (2021a) that these stringlets are the dominant contribution to the boson peak and grow upon heating, leading to the material softening upon heating. A similar phenomenon was originally observed in the glassy interfacial dynamics of crystalline Ni nanoparticles in their “premelted” state, where anharmonic interactions in the crystalline material are prevalent Zhang and Douglas (2013a, b). The existence of these structures and their significance for the boson peak have been confirmed in both 2D and 3D amorphous systems by Hu and Tanaka Hu and Tanaka (2022, 2023). Moreover, Betancourt et al. Betancourt P., Starr, and Douglas (2018); Betancourt et al. (2015) have shown a direct relation between the stringlets and string-like collective motion that has often been observed in connection with activated transport occurring on a much longer timescale. A general relation between the stringlet size and the mean square displacement on the fast $\beta$-relaxation time, a quantity that has been related to the long time $\alpha$-relaxation time, has been observed as well Douglas et al. (2016). Recently, Hu and Tanaka confirmed and extended these relations between the fast dynamics and $\alpha$-relaxation Hu and Tanaka (2022, 2023). Experimental studies have long suggested that transverse modes might give a predominant contribution to the boson peak Winterling (1975); Scopigno et al. (2003); Tian et al. (2021) and later simulation studies have provided strong support for this proposition Schober (2004); Shintani and Tanaka (2008). Additionally, it has also been noted that excess modes of a one-dimensional nature Novikov (1990) are consistent the commonly observed near linear frequency dependence of the light-vibrational coupling parameter $C(\omega)$ Achibat, Boukenter, and Duval (1993); Sokolov et al. (1993); Surovtsev et al. (2004). These observations and their interpretation are broadly in accord with the idea that topological “defects” can give rise to the BP in glasses Angell (2004); Bhat et al. (2006), and also with the related suggestion that these excitations are important for the plasticity and yield strength of glassy materials Baggioli et al. (2021); Baggioli (2023); Baggioli, Landry, and Zaccone (2022); Wu et al. (2023).
@@
Figure 1: Left panel: A concrete visualization of string-like (“stringlets”) excitations involving cooperative particle exchange motion, and corresponding to localized modes having a frequency around the BP value. Figure taken with permission from Hu and Tanaka (2023). Right panel: A stringlet atomic configuration of the Al-Sm metallic glass system Zhang et al. (2021a) at $550$K.
@@
We note that previous models Malinovsky, Novikov, and Sokolov (1991); Hu and Tanaka (2022, 2023) phenomenologically assumed that the excess modes responsible for the boson peak involve structures with a log-normal size distribution, but the physical basis of this assumption has been completely unclear. Lund Lund (2015) recently sought to describe these excess normal modes using a continuum solid model in which a random distribution of line “defects” or “elastic strings” was assumed to exist based on simulation observations. Lund’s model is an extension of the elastic string model of Granato and Lücke Granato and Lücke (1956), introduced long ago in the description of the plasticity in highly defective crystals, where the strings correspond to pinned segments of dislocation lines in an extended dislocation network. We note that an even earlier string model of the boson peak was proposed by Novikov Novikov (1990). Lund’s model Lund (2015) was able to rationalize many of the anomalous features of glasses, including the presence of a BP anomaly, Rayleigh sound attenuation constant, a negative sound dispersion for acoustic phonons with a minimum near the BP, and the saturation of the IR limit for transverse phonons. Nevertheless, the application of this model Bianchi, Giordano, and Lund (2020) to estimate the length distribution $p(l)$ of these hypothetical string-like structures, based on experimental observations of the VDOS in glycerol and silica, unfortunately led to a string length distribution inconsistent with simulation observations Zhang et al. (2021a); Hu and Tanaka (2022, 2023), which appears to indicate an approximate universal exponential form, $p(l)\approx\exp(-l/\lambda)$. We address this inconsistency below. Before continuing, we emphasize that despite several commonalities, stringlets are strictly speaking not “structural defects” as found in crystalline materials, even if they share some geometric similarities to dislocation networks. In this context, “string-like” dislocation segments exist within the dislocation loop networks where the points of impingement of the dislocation loops define pinning points or end-points of the dislocation segments Granato and Lücke (1956); Maurel et al. (2005a). As we shall discuss extensively below, string-like structures generally arise also in glass-forming liquids, which apparently have the same length distribution assumed in the modeling of dislocation segments. This situation allows us to model these structures based on a similar coarse-grained string mathematical model. We emphasize, however, that the stringlets of glass-forming liquids should not be equated with dislocation segments arising in mechanically deformed crystalline materials. A previous discussion of this important point can be found in section V of Ref.Lund (2015), and we discuss this “analogy” further in Appendix A. Moreover, this form of collective motion on the timescale of the fast beta relaxation process, occurring generally on a timescale on order of a ps, should not be identified with collective motion associated with thermal activation events associated with the alpha relaxation process that occurs on a much longer timescale associated with the average rate of intermittent hopping events associated with particle diffusion and the structural relaxation of the material as a whole Betancourt P., Starr, and Douglas (2018); Pazmiño Betancourt, Douglas, and Starr (2014). As we shall discuss below, Hong et al. Hong, Novikov, and Sokolov (2011a, b) have heuristically identified the boson peak with this alternative form of collective motion.
@@
Here, we revisit the Lund string model of the boson peak Lund (2015) and extend it to make a parameter-free prediction of $\omega_{BP}$ in both 2D and 3D for amorphous solids. In contrast to the work of Lund, however, we assume an exponential form for the string length distribution $p(l)$, at the outset to be consistent with simulation observations Zhang et al. (2021a); Hu and Tanaka (2022, 2023), and we further assume that the average stringlet vibration speed equals the transverse sound velocity to uniquely fix our model parameters. This model leads an extremely simple analytic expression for the BP frequency without any fitting parameters, which appears to be highly predictive in the the low temperature glass state, where many boson peak measurements are normally performed. In particular, our predictions are in good agreement with the experimental results for the BP frequency at $T=0$, with an uncertainty between $1\,\%$ and $5\,\%$. The stringlet model also captures the correct qualitative trend at high temperatures where these vibrational modes begin to become overdamped. Interestingly, we are also able to reproduce the softening of the BP frequency with temperature and to explain the experimentally observed scaling of $\omega_{BP}$ with the shear modulus $G$ Tomoshige et al. (2019). Our theory suggests that a strong damping of the stringlet modes is required above $T_{g}$ to describe simulation observations. As a by-product, the introduction of a friction term for the string vibrations produces a low-frequency divergence in the VDOS normalized to the Debye law, $g(\omega)/\omega^{d-1}$, which is ubiquitously observed in experimental and simulation data at finite temperature Hong, Novikov, and Sokolov (2011a); Buchenau (1993) (resembling similar features in anharmonic crystals and incommensurate structures), and usually not discussed in most of the traditional theories for the BP (see Cano and Levanyuk (2004); Baggioli and Zaccone (2019b, c, 2021); Jiang et al. (2023) for some exceptions).
@@ II The stringlet model
We initially directly follow Lund et al. Lund (2015); Bianchi, Giordano, and Lund (2020), and model the stringlets as 1D vibrating strings that are anchored at both ends. Similar results follow by assuming that the strings are free at both ends, since, in both cases, the vibrations would be characterized by a fundamental wave-vector $k=\pi/l$, with $l$ the length of the string. As noted before, 1D vibrational modes being excited by light are consistent with the low-frequency Raman scattering observations on glasses, where a near linear scaling for the light to vibration coupling coefficient $C(\omega)$ has been found over a large frequency range Novikov and Surovtsev (1999). According to this simple continuum based model, which is agnostic to the actual physical origin of these vibrational modes, each stringlet of length $l$ has a corresponding frequency of vibration given by,
@@
where $v$ is the speed of propagation of the stringlet wave, and $\gamma$ represents a “friction parameter” (see Lund (2015); Bianchi, Giordano, and Lund (2020) for details). Eq.(1) is based on the assumption that only the fundamental frequency of vibration of the stringlet is relevant, i.e., higher harmonics are neglected. We consider a set of stringlets whose length follows a distribution denoted as $p(l)$. Based on this assumption, the stringlet density of states $g_{s}(\omega)$ can readily be obtained using the inverse distribution function. More precisely, given the distribution function of a random variable $X>0$, the distribution function of the inverse variable $Y\equiv 1/X$ can be derived by relating the cumulative distribution function of the initial variable $F(x)$ to that of the inverse variable,
@@
where $d$ is the spatial dimension. Notice how the minus sign in the above equation, which can be formally derived using the procedure described above, ensuring the physically required positivity of the stringlet density of states $g_{s}(\omega)$ presented below. In three dimensions ($d=3$), the factor of $d-1=2$ in the r.h.s. of Eq.(4) reflects the ability of the stringlets to oscillate along two linearly independent directions that are perpendicular to its equilibrium orientation. Formally, this framework is equivalent to considering a distribution of damped oscillators with frequency $\omega_{0}$ and damping rate $\gamma$. We can then define the quality factor $Q=\omega_{0}/\gamma$ according to which $Q\ll 1$ corresponds to underdamped vibrations and $Q$ of $\mathcal{O}(1)$ to overdamped ones. Notice that a priori $Q$ depends on the stringlet length $l$ since $\omega_{0}$ exhibits this property.
We next introduce a new assumption – with respect to previous work Lund (2015); Bianchi, Giordano, and Lund (2020) – into this type of quasi-continuum framework that is an exponential distribution for the stringlet length,
@@
which is motivated directly by simulation observations Zhang et al. (2021a); Hu and Tanaka (2022, 2023). Here, $\lambda$ is the average stringlet length. Notice that this distribution is different from what assumed in Bianchi, Giordano, and Lund (2020), where a localized Gaussian-like distribution of lengths is considered (see Fig.7 therein). As shown in Zhang et al. (2021a); Hu and Tanaka (2022, 2023), this assumption is inconsistent with the simulation data. From simple physical arguments, it is clear that the stringlet length $l$ has to be larger than the atomic scale $a$ and smaller than the sample size $L$. For this reason, Eq.(5) represents the correct physical distribution in the regime $a\ll l\ll L$. Importantly, the average stringlet length $\lambda$, that as we will see corresponds to the frequency range around the boson peak, is of the order of nanometers. That is at least one order of magnitude larger than $a$ and much smaller than $L$. This implies that, in first approximation, any correction to (5), including the existence of UV cutoff set by $a$, produces subleading effects to the density of states around the BP frequency, which is the main subject of this work. In other words, for frequencies of the order of the BP scale, Eq.(5) is the correct distribution function.
@@
In the limit of zero damping, well below $T_{g}$, this expression becomes particularly simple,
$\lim_{\gamma\rightarrow 0}g_{s}(\omega)=p_{0}\,\pi(d-1)\,v\,\frac{\,e^{-\frac{\pi v}{\lambda\,\omega}}}{\omega^{2}}\,.$ (7)
We repeat that the above expression is not expected to capture the correct physics for very large frequencies, of the order of $\omega_{\text{UV}}\approx v/a$, with $a$ the atomic distance. The main reason is that stringlet vibrations cease to exist at that microscopic scale and the stringlet density of states has to be corrected with a cutoff similar to the Debye frequency for acoustic phonons. However, such a microscopic frequency scale is much larger than the BP frequency and therefore irrelevant for the present discussion. From Eq.(6), we can see that in two spatial dimensions the Debye normalized stringlet VDOS $g_{s}(\omega)/\omega$ has a maximum $\omega_{max}$,
@@
More generally, as discussed previously in Pazmiño Betancourt, Douglas, and Starr (2014), and below, the stringlet distribution is not a perfect exponential, as expected from mean field theories of linear string formation at equilibrium. A better, more general, representation involves a combination of an exponential form with a power-law prefactor Pazmiño Betancourt, Douglas, and Starr (2014),
@@
where the exponent $\theta$ depends on the topology of the string-like structures, and their excluded volume and other interactions. It is straightforward to re-derive the stringlet VDOS using this distribution. This yields the relation,
$\tilde{g}_{s}(\omega)=\frac{(d-1)\pi^{1-\theta}p_{0}\,\omega\left(\frac{v}{\sqrt{\gamma^{2}+\omega^{2}}}\right)^{3-\theta}e^{-\frac{\pi v}{\lambda\sqrt{\gamma^{2}+\omega^{2}}}}}{v^{2}}$ (12)
@@
$\lim_{\gamma\rightarrow 0}\tilde{g}_{s}(\omega)=(d-1)\pi^{1-\theta}p_{0}\,v^{1-\theta}\omega^{\theta-2}e^{-\frac{\pi v}{\lambda\omega}}.$ (13)
@@
Finally, we consider the full vibrational density of states $g(\omega)$, including the phononic Debye contribution. We then arrive at the simple relation,
$g(\omega)=g_{\text{Debye}}(\omega)+g_{s}(\omega),$ (15)
@@
where $v_{L},v_{T}$ are respectively the speed of longitudinal and tranverse acoustic phonons. The arguments above indicate that the boson peak frequency $\omega_{BP}$ coincides with the location of the maximum in the stringlet contribution to the VDOS, $g_{s}(\omega)$. The independence of the two contributions to the total density of states in Eq.(15) relies on the simplifying assumption that phonons and stringlets are non-interacting. As we will see below, this simple expression is sufficient to capture the most salient features related to the boson peak. It is certainly possible to go beyond this leading order approximation by considering the interactions between these no longer independent normal modes. Formally, this might be done by modifying the respective dynamical equations, in a way similar to what done for dislocations and phonons in ordered crystals (see for example Maurel et al. (2005a, b)). We leave this refinement of the model to future work.

### v3

@@ Abstract
The boson peak (BP), a low-energy excess in the vibrational density of states over the Debye contribution, is often identified as a characteristic of amorphous solid materials. Despite decades of efforts, its microscopic origin still remains a mystery. Recently, it has been proposed, and corroborated with simulations, that the BP might stem from intrinsic localized modes involving one-dimensional (1D) string-like excitations (“stringlets”). We build on a theory originally proposed by Lund that describes the localized modes as 1D vibrating strings, but we specify the stringlet size distribution to be exponential, as observed in simulations. We provide an analytical prediction for the BP frequency $\omega_{BP}$ in the temperature regime well below the observed glass transition temperature $T_{g}$. The prediction involves no free parameters and accords quantitatively with prior simulation observations in 2D and 3D model glasses based on inverse power law potentials. The comparison of the string model to observations is more uncertain when compared to simulations of an Al-Sm metallic glass material at temperatures well above $T_{g}$. Nonetheless, our stringlet model of the BP naturally reproduces the softening of the BP frequency upon heating and offers an analytical explanation for the experimentally observed scaling with the shear modulus in the glass state and changes in this scaling in simulations of glass-forming liquids.
Finally, the theoretical analysis highlights the existence of a strong damping for the stringlet modes above $T_{g}$, which leads to a large low-frequency contribution to the 3D vibrational density of states, observed in both experiments and simulations.
@@ I Introduction
Amorphous solids exhibit a variety of characteristic anomalies in their vibrational, thermodynamic and transport properties, when compared to the more familiar case of ordered crystalline matter [1; 2]. Among the different, and apparently universal [3], anomalous properties of glassy materials, the boson peak (BP) is probably the most commonly discussed and controversial. This peak refers to an excess in the vibrational density of states (VDOS) normalized by the phonon density of states of Debye theory, $g(\omega)/\omega^{d-1}$, where $d$ is the number of spatial dimensions and $\omega$ the frequency. The BP is ubiquitously observed in amorphous systems [4] and recently this feature has also been observed in crystalline systems at finite temperature (e.g., [5; 6]), so it is unclear whether it arises from structural disorder [7].
@@
Several theoretical models have been proposed to rationalize the BP anomaly in glasses [8; 9; 10; 11; 12; 13; 14; 15; 16; 7; 17; 18; 19; 20; 21], but a final verdict has not yet been reached. A common idea is that the BP represents a signature for the emergence of additional vibrational modes coexisting, but distinct from phonons. These “excess modes” have been suggested to arise from both structural disorder or from the anharmonicity in intermolecular interactions inherent to the liquid state and heated crystals. These commonly observed vibrational modes are expected to be effectively localized as they involve a rather limited number of atoms or molecular segments in complex molecules [14; 15; 22; 23; 24; 25; 26; 27; 28]. At large enough wave-vector, they also interact with phonons, making them overdamped, i.e., reaching the Ioffe-Regel (IR) limit [29; 30; 31], leading to hindered thermal transport because of the consequent strong phonon scattering.
@@
Despite numerous experimental and simulation studies, a fully predictive theoretical model explaining the existence of these modes and how they give rise to the observed boson peak remains elusive. Recently, it has been recognized that, on atomic scales, localized string-like excitations that involve reversible particle exchange motion in the form of linear polymeric structures exist in glass-forming liquids. These dynamic structures have been termed “stringlets[32]. See Fig.1 for a visual representation from simulations. Moreover, it has been shown [33; 32] that these stringlets are the dominant contribution to the boson peak and grow upon heating, leading to the material softening upon heating. A similar phenomenon was originally observed in the glassy interfacial dynamics of crystalline Ni nanoparticles in their “premelted” state, where anharmonic interactions in the crystalline material are prevalent [33; 34]. The existence of these structures and their significance for the boson peak have been confirmed in both 2D and 3D amorphous systems by Hu and Tanaka [35; 36]. Moreover, Betancourt et al. [37; 38] have shown a direct relation between the stringlets and string-like collective motion that has often been observed in connection with activated transport occurring on a much longer timescale. A general relation between the stringlet size and the mean square displacement on the fast $\beta$-relaxation time, a quantity that has been related to the long time $\alpha$-relaxation time, has been observed as well [39]. Recently, Hu and Tanaka confirmed and extended these relations between the fast dynamics and $\alpha$-relaxation [35; 36]. Experimental studies have long suggested that transverse modes might give a predominant contribution to the boson peak [40; 41; 42] and later simulation studies have provided strong support for this proposition [43; 29]. Additionally, it has also been noted that excess modes of a one-dimensional nature [44] are consistent the commonly observed near linear frequency dependence of the light-vibrational coupling parameter $C(\omega)$ [45; 46; 47]. These observations and their interpretation are broadly in accord with the idea that topological “defects” can give rise to the BP in glasses [48; 49], and also with the related suggestion that these excitations are important for the plasticity and yield strength of glassy materials [50; 51; 52; 53].
@@
Figure 1: Left panel: A concrete visualization of string-like (“stringlets”) excitations involving cooperative particle exchange motion, and corresponding to localized modes having a frequency around the BP value. Figure taken with permission from [36]. Right panel: A stringlet atomic configuration from the simulation Al-Sm metallic glass system [32] at $550$K.
@@
We note that previous models [54; 35; 36] phenomenologically assumed that the excess modes responsible for the boson peak involve structures with a log-normal size distribution, but the physical basis of this assumption has been completely unclear. Lund [55] recently sought to describe these excess normal modes using a continuum solid model in which a random distribution of line “defects” or “elastic strings” was assumed to exist based on simulation observations. Lund’s model is an extension of the elastic string model of Granato and Lücke [56], introduced long ago in the description of the plasticity in highly defective crystals, where the strings correspond to pinned segments of dislocation lines in an extended dislocation network. We note that an even earlier string model of the boson peak was proposed by Novikov [44]. Lund’s model [55] was able to rationalize many of the anomalous features of glasses, including the presence of a BP anomaly, Rayleigh sound attenuation constant, a negative sound dispersion for acoustic phonons with a minimum near the BP, and the saturation of the IR limit for transverse phonons. Nevertheless, the application of this model [57] to estimate the length distribution $p(l)$ of these hypothetical string-like structures, based on experimental observations of the VDOS in glycerol and silica, unfortunately led to a string length distribution inconsistent with simulation observations [32; 35; 36], which appears to instead indicate an approximate exponential form, $p(l)\sim\exp(-l/\lambda)$. We address this inconsistency below. Before continuing, we emphasize that despite several commonalities, stringlets are strictly speaking not “structural defects” as found in crystalline materials, even if they share some geometric similarities to dislocation networks. In this context, “string-like” dislocation segments exist within the dislocation loop networks where the points of impingement of the dislocation loops define pinning points or end-points of the dislocation segments [56; 58]. As we shall discuss extensively below, string-like structures generally arise also in glass-forming liquids, which apparently have the same length distribution assumed in the modeling of dislocation segments. This situation allows us to model these structures based on a similar coarse-grained string mathematical model. We emphasize, however, that the stringlets of glass-forming liquids should not be equated with dislocation segments arising in mechanically deformed crystalline materials. A previous discussion of this important point can be found in Section V of Ref.[55], and we discuss this “analogy” further in Appendix A. Moreover, this form of collective motion on the timescale of the fast beta relaxation process, occurring generally on a timescale on order of a ps, should not be identified with collective motion associated with thermal activation events associated with the alpha relaxation process that occurs on a much longer timescale associated with the average rate of intermittent hopping events associated with particle diffusion and the structural relaxation of the material as a whole [37; 59]. As we shall discuss below, Hong et al. [60; 61] have heuristically identified the boson peak with this alternative form of collective motion.
@@
Here, we revisit the Lund string model of the boson peak [55] and extend it to make a parameter-free prediction of $\omega_{BP}$ in both 2D and 3D for amorphous solids. In contrast to the work of Lund, however, we assume an exponential form for the string length distribution $p(l)$, at the outset to be consistent with simulation observations [32; 35; 36], and we further assume that the average stringlet vibration speed equals the transverse sound velocity to uniquely fix our model parameters. This model leads an extremely simple analytic expression for the BP frequency without any fitting parameters, which appears to be highly predictive in the the low temperature glass state, where many boson peak measurements are normally performed. In particular, our predictions are in good agreement with simulation results for the BP frequency at $T=0$, with an uncertainty between $1\,\%$ and $5\,\%$. The stringlet model also captures the correct qualitative trend at high temperatures where these vibrational modes begin to become overdamped. Interestingly, we are also able to reproduce the softening of the BP frequency with temperature and to explain the experimentally observed scaling of $\omega_{BP}$ with the shear modulus $G$ [62]. Our theoretical model of the boson peak suggests that a strong damping of the stringlet modes is required above $T_{g}$ to describe simulation observations. As a by-product, the introduction of a friction term for the string vibrations produces a low-frequency divergence in the VDOS normalized to the Debye law, $g(\omega)/\omega^{d-1}$, which is ubiquitously observed in experimental and simulation data at finite temperature [60; 63] (resembling similar features in anharmonic crystals and incommensurate structures), and usually not discussed in most of the traditional theories for the BP (see [64; 65; 66; 67; 68] for some exceptions).
@@ II The stringlet model
We initially directly follow Lund et al. [55; 57], and model the stringlets as 1D vibrating strings that are anchored at both ends. Similar results follow by assuming that the strings are free at both ends, since, in both cases, the vibrations would be characterized by a fundamental wave-vector $k=\pi/l$, with $l$ the length of the string. As noted before, 1D vibrational modes being excited by light are consistent with the low-frequency Raman scattering observations on glasses, where a near linear scaling for the light to vibration coupling coefficient $C(\omega)$ has been found over a large frequency range [69]. According to this simple continuum based model, which is agnostic to the actual physical origin of these vibrational modes, each stringlet of length $l$ has a corresponding frequency of vibration given by,
@@
where $v$ is the speed of propagation of the stringlet wave, and $\gamma$ represents a “friction parameter” (see [55; 57] for details). Eq.(1) is based on the assumption that only the fundamental frequency of vibration of the stringlet is relevant, i.e., higher harmonics are neglected. We consider a set of stringlets whose length follows a distribution denoted as $p(l)$. Based on this assumption, the stringlet density of states $g_{s}(\omega)$ can readily be obtained using the inverse distribution function. More precisely, given the distribution function of a random variable $X>0$, the distribution function of the inverse variable $Y\equiv 1/X$ can be derived by relating the cumulative distribution function of the initial variable $F(x)$ to that of the inverse variable,
@@
where $d$ is the spatial dimension. Notice how the minus sign in the above equation, which can be formally derived using the procedure described above, ensuring the physically required positivity of the stringlet density of states $g_{s}(\omega)$ presented below. In three dimensions ($d=3$), the factor of $d-1=2$ in the r.h.s. of Eq.(4) reflects the ability of the stringlets to oscillate along two linearly independent directions that are perpendicular to its equilibrium orientation. Formally, this framework is equivalent to considering a distribution of damped oscillators with frequency $\omega_{0}$ and stringlet damping rate $\gamma$. We can then define the quality factor $Q=\omega_{0}/\gamma$ according to which $Q\ll 1$ corresponds to underdamped vibrations and $Q$ of $\mathcal{O}(1)$ to overdamped ones. Notice that a priori $Q$ depends on the stringlet length $l$ since $\omega_{0}$ exhibits this property.
We next introduce a new assumption – with respect to previous work [55; 57] – into this type of quasi-continuum framework that is an exponential distribution for the stringlet length,
@@
which is motivated directly by simulation observations [32; 35; 36]. The prefactor in Eq. (5) is determined by the distribution normalization. Notice that this distribution is different from what assumed in Ref. [57], where a localized Gaussian-like distribution of lengths is considered (see Fig.7 therein). As shown in Refs. [32; 35; 36], this assumption is inconsistent with simulation data.
The above distribution, Eq.(5), presents a natural lower cutoff $l_{0}$ that corresponds to the size of the smallest particle in the system. In particular, this implies the consistency condition $l>l_{0}$, i.e., physically a stringlet has to be larger in size than the size of an individual particle. The average stringlet length $\langle l\rangle$ imay be defined as,
[…]
When expressed in units of the smaller particle diameter, as done below, this relation reduces to $\langle l\rangle=1+\lambda$.
@@
Because of our current lack of understanding of the normalization of the stringlet density of states, and thus the total number of stringlets, we add an unknown constant in front of the above expression that takes into account the normalization of the stringlet density of states. This constant can only be determined by means of a microscopic theory of the stringlets or can otherwise be taken as a fitting parameter from the data, as done in our case. We therefore find it convenient to rewrite the above expression as,
$g_{s}(\omega)=\mathcal{N}_{s}\,\frac{\omega\,e^{-\frac{\pi v}{\lambda\,\sqrt{\gamma^{2}+\omega^{2}}}}}{\left(\gamma^{2}+\omega^{2}\right)^{3/2}}\,,$ (8)
where all the prefactors have been absorbed into the prefactor $\mathcal{N}_{s}$, an empirical parameter since its determination requires knowledge of the total number of stringlets. Before continuing, we notice that the value of $\mathcal{N}_{s}$ does not affect the prediction of the BP frequency but this quantity does influence its intensity. As one might expect, the stringlet model of the boson peak predicts that the larger the number of stringlets the stronger the intensity of the BP. We note that simulation could provide information about the number of stringlets, but no information of this kind has so far been collected in simulation studies. It would be particularly interesting to determine also if the number of stringlets correlates with the fragility of glass-formation.
In the limit of zero stringlet damping, a situation that applies physically well below $T_{g}$, Eq. (8) reduces to the particularly simple form,
$\lim_{\gamma\rightarrow 0}g_{s}(\omega)=\mathcal{N}_{s}\,\frac{\,e^{-\frac{\pi v}{\lambda\,\omega}}}{\omega^{2}}\,.$ (9)
From Eq.(8), we can see that in two spatial dimensions the Debye normalized stringlet VDOS $g_{s}(\omega)/\omega$ has a maximum $\omega_{max}$,
@@
More generally, as discussed previously in [59], and below, the stringlet distribution is not a perfect exponential, as expected from mean field theories of linear string formation at equilibrium. A better, more general, representation involves a combination of an exponential form with a power-law prefactor [59],
@@
where the exponent $\theta$ depends on the topology of the string-like structures, and their excluded volume and other interactions. It is straightforward to re-derive the stringlet VDOS using this distribution. Again, we absorb constants into a modified prefactor, $\mathcal{\tilde{N}}_{s}$. This yields the relation,
$\tilde{g}_{s}(\omega)=\mathcal{\tilde{N}}_{s}\,\omega\left(\gamma^{2}+\omega^{2}\right)^{(\theta-3)/2}e^{-\frac{\pi v}{\lambda\sqrt{\gamma^{2}+\omega^{2}}}}$ (14)
@@
$\lim_{\gamma\rightarrow 0}\tilde{g}_{s}(\omega)=\mathcal{\tilde{N}}_{s}\,\omega^{\theta-2}e^{-\frac{\pi v}{\lambda\omega}}.$ (15)
@@
Finally, we consider the full vibrational density of states $g_{\mathrm{total}}(\omega)$, including the phononic Debye contribution. We then arrive at the simple relation,
$g_{\mathrm{total}}(\omega)=g_{\text{Debye}}(\omega)+g_{s}(\omega),$ (17)
@@
where $v_{L},v_{T}$ are respectively the speed of longitudinal and tranverse acoustic phonons. The arguments above indicate that the boson peak frequency $\omega_{BP}$ coincides with the location of the maximum in the stringlet contribution to the VDOS, $g_{s}(\omega)$. The independence of the two contributions to the total density of states in Eq.(17) relies on the simplifying assumption that phonons and stringlets are non-interacting. As we will see below, this simple expression is sufficient to capture the most salient features related to the boson peak. It is certainly possible to go beyond this leading order approximation by considering the interactions between stringlets and phonons. Formally, this might be done by modifying the respective dynamical equations, in a way similar to what done for dislocations and phonons in ordered crystals (e.g., see Refs. [58; 70]). While this type of coupling is naturally expected to exist, our results below indicate that, at least in the low temperature limit, it is not fundamental in deriving the position of the BP frequency. Nevertheless, we expect such coupling to be crucial to describe other properties in amorphous solids such as sound attenuation and thermal conductivity.